OpenStudy (anonymous):

Evaluate x3 - 3y when x = -3 and y = 4.

OpenStudy (zpupster):

(-3)^3 - 3(4) =?

OpenStudy (gorv):

plug value of x and y (-3)^3-3*4= and (-3)^3=(-3)*(-3)*(-3)